The debt-equity ratio is 60%. The after-tax cost of debt is 15%. The cost of equity is 20%. What is the WACC? 18.1% 17.0% 18.0% 16.9%
Answer #1

WACC = Weight of Debt * After tax cost of debt + Weight of Equity * Cost of Equity

Debt-Equity Ratio = 60%

This implies if debt = 60, Equity = 100

Weight of Debt = Debt/(Debt + Equity) = 60/(60 + 100) = 37.50%

Weight of Equity = Equity/(Debt + Equity) = 100/(60 + 100) = 62.50%

WACC = 37.50% * 15% + 62.50% * 20%

WACC = 18.1%